OpenStudy (anonymous):

distance between (3,-1) (-2,-4)

OpenStudy (anonymous):

\[d^2=(x2-x1)^2 + (y2-y1)^2\]

OpenStudy (anonymous):

susbstitue values and square root it

OpenStudy (anonymous):

i did it and got 34 squared i need help on that

OpenStudy (anonymous):

\[d^2=34\]

OpenStudy (anonymous):

\[d = \sqrt{34}\]

OpenStudy (anonymous):

is it 4 and 8.5 squared

OpenStudy (anonymous):

it should be the square root of 9 + 25